OpenStudy (anonymous):

Solve P = 2(l + w) for l please

jimthompson5910 (jim_thompson5910):

P = 2(l + w) P = 2l +2w P - 2w = 2l (P - 2w)/2 = l l = (P - 2w)/2

OpenStudy (anonymous):

Thank you!